### v2

# Jets-separation thresholds, Seshadri constants and higher Gauss-Wahl maps on abelian varieties Nelson Alvarado N. Alvarado

Abstract.
Given a closed subscheme of a polarized abelian variety we define its vanishing threshold with respect to and relate it to the Seshadri constant of the ideal defining As a particular case, we introduce the notion of jets-separation threshold, which naturally arises as the vanishing threshold of the -infinitesimal neighborhood of a point. Afterwards, by means of Fourier-Mukai methods we relate the jets-separation thresholds with the surjectivity of certain higher Gauss-Wahl maps. As a consequence we obtain a criterion for the surjectivity of those maps in terms of the Seshadri constant of the polarization
In this article we introduce the vanishing threshold of a closed subscheme of a polarized abelian variety defined over an algebraically closed field. In the case of a closed point (with its reduced structure) this number is no other than the basepoint-freeness threshold introduced by Jiang-Pareschi in [cohrank], which has also been studied by Caucci [caucci], Ito [itobir] and Jiang [jiang], using methods from birational geometry, and by Rojas [Rojas], by means of stability conditions in the two-dimensional case. When we consider the -infinitesimal neighborhood of a closed point (i.e the scheme defined by the -power of the ideal of the point) a natural notion of -jets separation threshold arises.
In the first part of this work, we introduce the aforementioned thresholds and show some examples. Afterwards, we prove that if is a closed subscheme and is its -infinitesimal neighborhood then the sequence, suitably normalized, of vanishing thresholds of with respect to a polarization converges to the Seshadri constant with respect to of the ideal defining As a particular case, when is a point (with its reduced structure) we get that the sequence of -jets separation thresholds (suitably normalized) converges to the classical Seshadri constant of This is no surprise since, by a theorem of Demailly ([Dem, Theorem 6.4]), the Seshadri constant of a line bundle at a point reflects the asymptotic behaviour of jet separation. However, unlike the invariants considered by Demailly (see [PositivityI, Proposition 5.1.6] for the definition of these numbers), which are integers and depend on , the -jets separation thresholds are real numbers and do not depend on an specific closed point In this context, our asymptotic description can be seen as a finer version, valid for abelian varieties, of Demailly’s result.
Subsequently, we show that -jets separation thresholds are intimately related, via the Fourier-Mukai transform, to certain higher Gauss-Wahl maps. In fact, we prove that they carry esentially the same information of certain (suitably defined) -Gauss-Wahl surjectivity threshold. This paralells the fact, shown by Jiang-Pareschi in [cohrank, Proposition 8.1], that basepoint freeness threshold carries the same information as the surjectivity threshold for multiplication maps of global sections (and therefore, is intimately related to projective normality). As a byproduct of the aforementioned relation we obtain a description of the Seshadri constant in terms of the asymptotic behaviour of the surjectivity of certain Gauss-Wahl maps. Finally, by means of Nadel’s vanishing theorem, we bound the -jets separation thresholds in terms of the Seshadri constant, when we work in characteristic zero. As a consequence, we establish precise lower bounds for the Seshadri constant guaranteeing the surjectivity of higher Gauss-Wahl maps, asymptotically improving the results present in the literature.
Turning to a more detailed presentation, we now proceed to define the afore mentioned thresholds. First, we recall that a coherent sheaf on an abelian variety is said to satisfy the index theorem with index zero (IT(0) for short) if
where is the line bundle on parametrized by via a normalized Poincaré bundle.
Now, let be a polarization on (that is, an ample class in ). As shown by Jiang-Pareschi, the IT(0) condition can be extended for -twisted sheaves (we recall this notion in Section 2.1). In this context, we define the vanishing threshold of with respect to as the real number
In this context, we define the vanishing threshold of a closed subscheme of with respect to to be the vanishing threshold of its ideal. In symbols:
As an example, fix a closed point with ideal sheaf and consider a positive integer Let be a line bundle representing As is ample, all the line bundles are isomorphic to for some closed point where is the translation map, and hence the vanishing
means that for every the line bundle separates -jets at and thus, that separates -jets at every point of As the higher cohomology groups of authomatically vanish, the condition (1.2) turns to be equivalent to IT(0) and therefore, it is natural to define the -jets separation threshold of using the vanishing threshold with respect to of the -infinitesimal neighborhood of , where is a (any) closed point of that is:
Our first result relates the vanishing threshold of infinitesimal neighborhoods of a closed subscheme with the Seshadri constant of the ideal defining (see section 4 for its definition). More precisely, we prove the following:
Theorem 1.1.
Let be a -dimensional abelian variety and a closed subscheme with ideal For write for the closed subscheme of defined by the ideal Let be an ample line bundle on with class then:
The following inequalities hold:
If then
For with we have
The sequence converges and
If the base field has characteristic zero, the following inequalities of jets separation thresholds hold:
where is the Seshadri constant of
Our second result concerns higher Gauss-Wahl maps. They were introduced by Wahl in [Wahl] and they are a hierarchy of linear maps associated to a pair of line bundles , where is no other than the multiplication map of global sections. For the moment, we give an informal introduction to these maps (in Section 5 we recall their precise definition). In brief, given two line bundles on a projective variety the associated -th Gauss-Wahl map is a linear map
where is a sheaf that in the case that is very ample is no other than the -order conormal bundle associated to the embedding of into (for instance, can be identified with ).
When separates -jets at every point, the surjectivity of is detected by the vanishing of (we refer to section 5 for details). In this context, informally speaking, the vanishing threshold (see (1.1)) of the sheaf with respect to may be seen as the minimal “rational power” of such that the “rational” Gauss-Wahl map is surjective. Extending [cohrank, Proposition 8.1], our second result says that the numbers
are related, via the Fourier-Mukai transform (reviewed in Section 2.3), with the jets-separation thresholds as follows:
Theorem 1.2.
Let be an ample line bundle on an abelian variety Let be the class of and suppose that separates -jets at every point of (thus ). Then the following equality holds:
As a consequence, this gives a surprising expression of the Seshadri constant in terms of the asymptotic surjectivity of higher Gauss-Wahl maps. Concretely, we have the following:
Corollary 1.3.
Let be a polarization on an abelian variety Then:
An interesting consequence of this relation is the fact that we are able to characterize special kinds of polarized abelian varieties using Gauss-Wahl maps. Concretely, here we highlight two special cases:
1. Decomposable abelian varieties with a curve factor: In [Nakamaye] Nakamaye proved that if is a polarized abelian variety, then and, moreover, if and only if
where is a principally polarized elliptic curve and is another polarized abelian variety. In this context, Corollary 1.3 says that varieties of the form (1.3) are also characterized by the following equivalent conditions:
a)
for every
b)
the sequence is unbounded
2. Hyperelliptic jacobians: In [Deb, Theorem 7] it is shown that if is a curve of genus and is not hyperelliptic then while in the hyperelliptic case. Moreover, in the cited reference it is conjectured that the condition should characterize jacobians of hyperelliptic curves among indecomposable principally polarized abelian varieties (i.p.p.a.v) of dimension
In this context, Corollary 1.3 says that jacobians of hyperelliptic curves are characterized among jacobians (and conjecturally among all i.p.p.a.v) by the fact that for that is, for the failure of surjectivity of the Gauss-Wahl map for
Finally, as an application, we use Theorem 1.2 above to study the surjectivity of Gauss-Wahl maps where and are ample and algebraically equivalent line bundles. Our result is the following:
Theorem 1.4.
Let ample and algebraically equivalent ample line bundles on an abelian variety Let be positive integers and write for the class of and in Assume that
for some Then the Gauss-Wahl map is surjective.
For instance, the above result says that if that is, if separates -jets at every point, then is surjective, while is already surjective whenever On the other hand, using Theorem 1.1 3) we obtain a condition ensuring the surjectivity of certain Gauss-Wahl maps in terms of the Seshadri constant, which (at least asymptotically) improves the results present in the literature ([mult, Theorem 2.2], for example). More precisely, we get the following:
Corollary 1.5.
Let be a polarization on an abelian variery. Consider a positive integer such that where is the Seshadri constant of Then
In particular, for ample line bundles representing the Gauss-Wahl map is surjective as soon as
This article structures as follows: in Section 2 we survey the notions of generic vanishing for -twisted objects and their relation with the Fourier-Mukai functor, whose definition we also recall. In Section 3 we formally introduce the notion of vanishing threshold of a closed subscheme of an abelian variety and, in particular, the -jets separation thresholds. In this context, we establish Theorem 1.1 2.a). In Section 4 we discuss Seshadri constants of ideal sheaves, and relate them with the vanishing threshold of the corresponding closed subscheme, proving Theorem 1.1.1) and 2.b). In Section 5 we discuss Gauss-Wahl maps, introduce their surjectivity thresholds and prove a slightly more general version of Theorem 1.2 above by computing the Fourier transform of In Section 6, we prove Theorem 1.1 3) and show its consequences. In particular, in this section we prove Theorem 1.4 and Corollary 1.5. Finally, in Section 7 we state a couple of questions for future research: the first one regards the possibility to render our asymptotic results more effective and the second one regards the possibility to extend Theorem 1.1 3) to abelian varieties defined over fields of positive characteristic.
2. Preliminaries
In this section we recall the preliminary materials that we will need. We start by fixing some notation. Let be an abelian variety defined over an algebraically closed field . By a polarization on we mean an ample class For as before, we write for the dual of that is, is the abelian variety parametrizing traslation-invariant line bundles on For a polarization we write (or ) for the isogeny which sends a point to the point of corresponding to the line bundle where is the traslation map given by and is a (any) line bundle representing . For a positive integer we write for the isogeny Finally, we fix a normalized Poincaré bundle on and for write for the line bundle on parametrized by (equivalently, is the fiber of over ).
2.1. -twisted objects
For as before, we write for the bounded derived category of coherent sheaves on We consider pairs where is an object in is a rational number and is a polarization. On these pairs we impose the relation given by identifying with the pair for every integer and for every representing
Definition 2.1.
A -twisted object, written is the class of the pair under the relation described in the previous paragraph.
For a -twisted object we would like to define its cohomological support loci To start, let be a line bundle representing and consider with and Write
where stands for the sheaf (hyper)cohomology. These subsets depend on the choices of and For instance, if is another representant of then is just a traslation of On the other hand, if we change for with a positive integer not divisible by and we consider a symmetric , then:
where, to get the direct sum we used [Mumford, p.72] and the projection formula. However, we can make the following:
Definition 2.2.
For a -twisted object its -th cohomological support loci is the class of where is a representant of and are integers with and , under the equivalence relation generated by translations and by taking direct images of prime to the characteristic multiplication isogenies (however, see Section 2.2 below).
In particular, although is not a set, its dimension and codimension are well defined and they are denoted by and respectively. In [cohrank, Section 5] the cited authors use these loci in order to define generic vanishing notions for -twisted objects:
Definition 2.3.
We say that a -twisted object is
IT(0) if a (any) representant of is empty for
GV if for every
For example, if is a polarization, then from the index theorem (see [Mumford, III.16], or Kodaira vanishing, in the caracteristic zero case) it follows that is IT(0) for . Another important example of these generic vanishing conditions regards the notion of separation of jets that we now recall:
Definition 2.4.
Let be a variety and a closed point with ideal sheaf . Given an integer and a line bundle on we say that separates -jets at if the restriction map
is surjective.
In particular, separates -jets at if and only if is not a basepoint of Furthermore, if is not a basepoint of then separates 1-jets at if and only if separates tangent vectors at (i.e the map between tangent spaces induced by the rational morphism is injective).
Example 2.5.
Let be an abelian variety, be a closed point with ideal and an ample line bundle with class For a positive integer taking cohomology to the exact sequence we see that for every and every and thus for On the other hand, we see that a representant for is given by the set
Now, for an ample line bundle we have that for every Using this, we see that, up to a traslation, the above set is equal to
It follows that is GV if and only if there exists such that separates -jets at and it is IT(0) if and only if this happens for all In particular, is always GV while it is IT(0) if and only if is basepoint free.
We have the following characterization of the behaviour of the generic vanishing conditions of for varying :
Theorem 2.6([cohrank], Theorem 5.2).
A -twisted object is GV if and only if is IT(0) for every
If is GV but not IT(0) then for every the -twisted object is not GV
is IT(0) if and only if is IT(0) for every small enough
There is also the following important result of preservation of vanishing:
Theorem 2.7([RegAVIII], Theorem 3.2 and Remark 3.3 ).
Let coherent sheaves on an abelian variety and let be a polarization. Write for the derived tensor product of and
Then if both and are GV then is GV.
The conditions defined in this subsection can also be reformulated in terms of the Fourier-Mukai transform with kernel which we briefly survey in the following subsection.
2.2. Fourier-Mukai transform
If we write for the product projections from then we can consider the Fourier-Mukai functor given by
which by [Mukai] is an equivalence of categories. For an object we will write for the -th cohomology sheaf of
From the cohomology and base change theorem [EGAIII-II, Théorème (7.7.5) II] we have that
In particular, we see that if is IT(0) then for every Moreover, we have the following equivalence ([PPo3, Lemma 3.6]):
On the other hand, if is a positive integer (not necessarily coprime to the characteristic of the base field), then by [Mukai, (3.4)] we have that
In particular, as is finite, we see that the following conditions are equivalent:
a)
is GV
b)
for every and a (any) representation
c)
for every and a (any) representation
Example 2.8.
Applying to the exact sequence we get the exact sequence
and the at the right is supported on the points where fails to separate -jets.
2.3. Cohomological rank functions
For varying the fact that whether or not, can be expressed using the cohomological rank functions introduced by Jiang-Pareschi in [cohrank]. In this subsection we recall their definition and fix some notation
Definition 2.9.
Let be a -dimensional abelian variety, an ample class and Given the -th cohomological rank function of with respect to is the function given by
where is a (any) line bundle representing
Remark 2.10.
The minimum involved in the definition is actually the generic value of for varying In particular, from Example 2.5 we see that is GV if and only if
One of the main properties of the cohomological rank functions is the following fundamental transformation formula with respect to the Fourier-Mukai functor:
Proposition 2.11([cohrank], Proposition 2.3).
For and as in Definition 2.9 above and the following equalities hold:
a)
b)
where stands for the derived dual
Moreover, in Theorem 3.2 from loc-cit it is also shown that cohomological rank functions can be extended to continous functions
3. Vanishing thresholds
In this section we introduce the notion of vanishing thresholds for coherent sheaves. Afterwards, we give some examples and at the end we prove Theorem 1.1 1) from the introduction.
Definition 3.1.
Let be an abelian variety and Let be a polarization on The vanishing threshold of with respect to is the real number
Given a closed subscheme of and a non-negative integer we write
where is the ideal sheaf of
The -jets separation threshold of is the number
where we consider the reduced scheme structure on the set
We note that, by definition, we have that Also, thanks to Theorem 2.6 we have the following equality:
Example 3.2.
We have that is the basepoint-freeness threshold considered in [cohrank, Section 8]. In particular and if and only if has base points. More generally, from Theorem 2.6 we see that
if and only if there exists a representant of such that separates -jets at 0. Arguing as in Example 2.5, we see that this is equivalent to say that for any representant of there exists a point such that separates -jets at such a point.
if and only if for every representant of separates -jets at 0. Equivalently, any representant of separates -jets at every point of
Example 3.3.
Consider a principal polarization and a symmetric divisor representing it. If is indecomposable, then fails to separate 1-jets (i.e tangent vectors) just in the 2-torsion points of That is, is GV but not IT(0) which, by Remark 2.10 and Theorem 2.6, means that From [cav, p.99], the same conclusion holds if is decomposable.
Example 3.4.
Let be a -dimensional principally polarized abelian variety and a length two closed subscheme supported in the origin. Again in this case, we have that
Let be a symmetric divisor representing and identify with via We claim that for while To see this, notice that and hence we have an exact sequence
As and we obtain that for and an exact sequence
It follows that Now, for this example, the value of does not matter (however, see [GV-min-GV, Corollary 4.3]). Let By the transformation formula (Proposition 2.11b)) and the degeneration of the spectral sequence computing the hypercohomology we have that
From our previous calculation we obtain then that
As clearly for we conclude that
Example 3.5.
Let be a smooth curve of genus and an Abel-Jacobi map. By [cohrank, Theorem 7.5 and Proposition 7.6] we know that for On the other hand, from the exact sequence
it follows that
This means that
and hence is not GV for every On the other hand, in [RegAVI, Proposition 4.4] it is shown that is GV and thus
Now, by [GV-min-GV, Theorem 6.1] this is esentially the only example of a curve inside an abelian variety whose GV-threshold is at most one. More precisely, if is a smooth curve that generates an abelian variety and is GV for a principal polarization then has minimal class and thus by Matsusaka’s criterion [Mats, Theorem 3] it follows that
A couple of formal properties are the following:
Proposition 3.6.
a)
If is a smooth closed subscheme of an abelian variety then
where is the conormal sheaf of in
b)
The sequence of jets-separation thresholds is unbounded and not decreasing
Proof.
a) We want to prove that is GV whenever both and are GV. As is smooth, we have that
and thus we get exact sequences
and thus the claim follows from [itobir, Lemma 2.8 (1)].
b) The fact that the sequence is not decreasing follows from the previous part considering (with reduced scheme structure). Now, the sequence is unbounded and for every and for every line bundle we have
It follows that, for any fixed there exists such that for every the restriction map
is not surjective and thus can not be GV. In other words, for any there exists such that and hence the sequence is unbounded.
We also have the following:
Theorem 3.7.
Let be a polarization on an abelian variety and a closed subscheme of dimension at most one. Then for positive integers with the following inequalities hold:
In particular, the sequence converges and
In order to prove the theorem, the following subadditivity lemma is fundamental:
Lemma 3.8.
Let ideals sheaves of subschemes of dimension at most one. Let be a line bundle such that is GV. Then both and are GV.
Proof of the Lemma:.
For an object we write In this context, the hypothesis means that
Now, we claim that
It is clear that, once the claim is proved, the result follows. Now, to prove the claim, we prove the opposite contentions between the complements, that is, we show that if for we have then and if such vanishing holds then To do this we use the (fourth-quadrant) spectral sequence ([Huybrechts, (3.5)])
Tensoring the exact sequence by we see that the support of the sheaves for is contained in the intersection of the cosupports of and As such scheme has dimension at most one, it follows that
and the spectral sequence already degenerates at the second page. It follows that is a subquotient of and thus
as we wanted to see.
Now, to prove that is also GV, we note that we have the exact sequence
and thus taking cohomology we get an exact sequence
As the at the right is zero for and hence we get the inclusion
that we wanted to prove.
Proof of the Theorem:.
Let such that both and are GV. By 2.7 we have that is GV. If we write and the latter condition means that is GV. Now, as pullbacks commutes with derived tensor products and is a flat morphism (so and are also ideals), Lemma 3.8 above implies then that is GV, as we wanted to see.
Finally, from Fekete’s subadditivity lemma (see, [mustata, Lemma 1.4], for example), it follows that the sequence converges and
An immediate consequence is the following:
Corollary 3.9.
For every the following inequalites of jets-separation thresholds hold:
and if then for every
Remark 3.10.
Note that using the above corollary we see that if then and thus separates -jets at every point. In particular we recover the well known fact ([Bauer, Theorem 1],[RegAVIII, Theorem (4)]) that for every we have that separates -jets at every point.
4. Seshadri constants
In Section 3 above we proved that for a closed subscheme of dimension at most one, the sequence converges. In this section we establish that the limit is actually the inverse of the Seshadri constant of the ideal with respect to .
First, recall that for an ample line bundle and an ideal sheaf the Seshadri constant of with respect to is the real number
where is the blow-up along with exceptional divisor
Theorem 4.1.
Let be an ample line bundle on an a -abelian variety and the corresponding polarization. Let be a closed subscheme of Then:
If then
In particular, in this case we have that for every
Proof.
The proof of this statement closely follows [ELC, Theorem 3.2]. First we prove that In order to do this we consider a positive rational number with (note that this kind of numbers form a dense subset of ) and we need to show that there exists (possibly very big) such that for which is enough to find such that
To do this, we start by noticing that by [ELC, Lemma 3.3] there exists such that
for all all and all line bundles where and is the corresponding projection. Now, let be an integer such that For all such consider and thus we have
where in the last step we used the fact that to get the direct sum ([Mumford, p.72]). So, in order to get the vanishing (4.1), we need to ensure the vanishing of each direct summand.
Now, by (4.2) and the way we choose we have that
Now, as we have that is ample and thus, as we just need to consider finite ’s, by Serre’s vanishing we can take such that the desired vanishings (4.1) hold. Summarizing, we have established that is IT(0) for and thus
As can be arbitrarily close to we conclude that
To prove the opposite inequality, we first note that by [PositivityI, Proposition 5.4.5] we have that
while Mumford’s theorem ([PositivityI, Theorem 1.8.5]) says that
where being -regular with respect to means that for . Now, by definition of and Theorem 2.6, we have that is IT(0) for every . In particular, for any integer we have the following vanishings:
For we automatically have and hence is -regular with respect to as soon as
for which it is enough to have In particular, we obtain that
It follows that
Passing to the limit we get
and therefore
as we wanted to see.
If Theorem 3.7 above implies that we have equalities.
5. Gauss-Wahl maps
In this section we recall the definition of the Gauss-Wahl maps and in particular we introduce some thresholds that may be thought as the surjectivity thresholds of such maps. Then we establish Theorem 1.2 from the introduction, which relates these thresholds with the jets-separation thresholds.
To start, consider a smooth projective variety We write for the projections, for the diagonal and for the corresponding ideal sheaf. For we write for the closed subscheme of defined by the ideal (we consider ) thus
Definition 5.1.
Let be a line bundle on Given the sheaf of -principal parts of over (or the sheaf of -jets of over ) is the sheaf
We also consider the sheaf of -jets-relations
From the definition it follows that and Now, applying to the exact sequence we get an exact sequence
where the fiber over of the last arrow is given by the restriction
In particular, the right arrow is surjective if and only if separates -jets at every point of In this way we also see that and is the kernel of the evaluation map (which in [cohrank] is denoted by ).
On the other hand, applying to the exact sequence
we get an exact sequence
fitting in the following commutative and exact diagram:
where the surjectivity at the bottom follows from the fact that the sheaf is supported on the diagonal (for example, see [Ricolfi, Proposition 1.3] for details). In particular, from the 5-lemma it follows that is surjective as soon as the evaluation map is surjective, that is, as soon as separates -jets at every point of
Definition 5.2.
Given line bundles on a smooth projective variety and
The -th multiplication map associated to is the map
From the diagram (5.3) we see that the surjectivity of implies the surjectivity of Moreover, from (5.2) and the remark below it, we see that if generate -jets at every point then the surjectivity of is equivalent to the vanishing of and hence such vanishing implies the surjectivity of In the case that and is an abelian variety, this suggests to study the cohomological rank function where is the class of in
Lemma 5.3.
Let be an ample line bundle on an abelian variety of dimension . Then:
a)
For we have that is IT(0) and
b)
If separates -jets at every point then we have for every and while
for with equality if
Proof.
a) Write with and Without loss of generality we may assume that is symmetric and hence Applying to the exact sequences
we get that for every we have
and thus, inductively, we see that
as we wanted to prove. It also follows, inductively, that
b) Now, write with a positive integer and . As separates -jets at every point, the sequence in (5.1) is also right exact. As for applying and taking cohomology we obtain then the following isomorphisms
where the last equality comes from part a). Moreover, if then and hence the above isomorphism also holds for as soon as Now, if then and therefore we have the following exact sequence:
where the group at the right is zero if The result then follows from part a).
In the spirit of [cohrank, Section 8] we introduce the following numbers:
Definition 5.4.
Let be an ample line bundle on an abelian variety For we write
where is the class of in
Note that, when separates -jets at every point, the above lemma tells us that is the vanishing threshold (see Definition 3.1) and, in this context, Note also that the invariant is already considered in [cohrank], where it is denoted by In the cited reference it is shown that there is a relation between such number and the base-point freeness threshold, when is globally generated. In the following we compute to show that there is an analagous relation between and
Before stating our result, we need to introduce further notation.
Let be an ample line bundle on an abelian variety and For and we write
and for we write
and it is easy to see that this number does not depend on the particular representation As observed before, when separates -jets at every point, is no other than
Theorem 5.5.
Let be an abelian variety. Let be an ample line bundle on with class Then for every the following equality holds:
In particular, if separates -jets at every point, the following equality holds:
and if and are algebraically equivalent then
The main tool to prove this is the following lemma:
Lemma 5.6.
Let be an abelian variety and an ample line bundle over We write for the closed subscheme of defined by the ideal Then we have
Proof.
Let be the multiplication map and be the projections.
By [Mukai, (3.10)] we have that
Now, let be the map given by whose inverse is given by where is the difference map. Then we have:
where in the last line we used the fact that and the definition of the natural map in (5.1).
Proof of the Theorem:.
We have that is isomorphic in to the complex (concentrated in degrees 0 and 1). By Lemma 5.6 above, this implies that in we have the following isomorphism
where is the complex (concentrated in degrees 0 and 1) given by
By the transformation formula (Proposition 2.11) it follows that for we have
Now, by Lemma 5.3 above, we have that both and are IT(0) for In this context, writing with the spectral sequence ([Huybrechts, Remark 2.67])
says that
and thus
Finally, as setting the result follows.
An important consequence of Theorem 5.5 is that we can ensure the surjectivity of certain Gauss-Wahl maps when we know that is small. We discuss this application in detail in the next section. For the moment, we limit ourselves to discuss formal consequences of the above theorem and its relation with the Seshadri constant.
First, we note that, unlike the jets-separation thresholds, a priori it is not clear how changes when we replace by a multiple of it. However, using the previous theorem, we are able to give such a formula:
Corollary 5.7.
Let be an ample line bundle and suppose that separates -jets at every point. Then for every we have that
Proof.
As separates -jets, it follows that also separates -jets for every positive and hence, applying the previous theorem we get that the following equalities hold:
Combining with the results from previous section we get an expression of the Seshadri constant in terms of the thresholds
Corollary 5.8.
Let be an ample line bundle on an abelian variety, then
From Remark 3.10 we see that if is an ample line bundle then separates -jets at every point. Therefore, by Theorem 5.5 we have
On the other hand, by definition we know that Therefore we have
In particular the limit exists as an element of and by Proposition 4.1 it follows that
or, in other words, we have
as we wanted to see.
Using the results of this section, we get the following:
Corollary 5.9.
Let be an abelian variety of dimension and a polarization Then the following are equivalent:
a)
b)
for every
c)
For a line bundle representing the sequence is unbounded
d)
There exists a principally polarized elliptic curve and a -dimensional polarized abelian variety such that
Proof.
The equivalence between a) and b) follows as a combination of Theorem 3.7 and Theorem 4.1. From Corollary 5.8 above, the equivalence a)c) is immediate. Finally, the equivalence a)d) is the content of Nakamaye’s theorem [Nakamaye, Theorem 1.1].
6. Effective surjectivity of Gauss-Wahl maps
In this section we establish Theorem 1.4 from the introduction.
Theorem 6.1.
Let and be ample and algebraically equivalent line bundles on an abelian variety Let and Then the Gauss-Wahl map
is surjective whenever
In particular, if separates -jets at every point then is surjective.
Proof.
The surjectivity of (6.1) follows as soon as Indeed, by Lemma 5.3b) this inequality means that the -twisted sheaf is IT(0), where is the class of and in which in particular means that and thus that is surjective.
Now, we note that by hypothesis we have
which means that that is, separates -jets at every point of From Theorem (5.5) it follows that
Now, as the function is strictly increasing, we have
and hence the result follows.
Regarding the last part, we note that hypothesis means that On the other hand, if then and thus the result follows.
Finally, we use multiplier ideals in order to bound the jets-separation thresholds in terms of the Seshadri constant of To do this, we will need the following:
Lemma 6.2.
Let be an abelian variety and an ample line bundle on it. Fix a rational number with Then for every with and for every divisible enough integer, there exists a divisor such that
i)
ii)
iii)
is a log-resolution of and the strict transform is reduced
Proof.
Up to minor modifications, this is the content of [LPP, Lemma 1.2].
Now, if we fix and as in the previous lemma then considering and with we obtain:
Now, as is a log-resolution of we have that is snc and hence, as it follows that
In this way we can compute the multiplier ideal
On the other hand:
which is ample because of the way in we chose Therefore, the following inequality directly follows from Nadel’s vanishing:
Proposition 6.3.
For a polarization on an abelian variety the following inequality holds
where
and is a line bundle representing
Now, from Proposition 6.3 and Theorem 6.1 we are able to prove the following:
Corollary 6.4.
Let be ample and algebraically equivalent line bundles and Consider a positive integer such that Then the Gauss-Wahl map is surjective as soon as
Proof.
From Theorem 6.1 and Proposition 6.3, the desired surjectivity follows as soon as
where is the class of and Now, as the above inequality becomes (6.2).
To illustrate the usage of this result, let be algebraically equivalent ample line bundles as in the theorem. Consider two functions and suppose that we want to ensure the surjectivity of for all and In order to apply the corollary we would just need
or, equivalently:
As an example, we contrast with [mult, Theorem 2.2]. In such reference it is proved that is surjective for all with Now, if we consider we get that is surjective for all (no matter ) whenever
As as from Nakamaye’s theorem we get that if is not as in (1.3) then (6.3) holds for In this sense, Theorem 6.1 above asymptotically improves [mult, Theorem 2.2].
7. Further questions
1. In Corollary 5.9 we saw that a special kind of polarized abelian varieties, namely the products
are characterized by the equality for all Now, by [Steffens, Proposition 2] we know that if is a principally polarized abelian surface (p.p.a.s) then whenever is not a product. From Proposition 6.3 this means that for indecomposable p.p.a.s and we have
and thus, a posteriori, the polarized surfaces of the form (7.1) are characterized by the equality
Question 7.1.
Let be a positive integer. Does there exist a positive integer such that a -dimensional principally polarized abelian variety is as in (7.1) if and only if ?
More generally, we may ask:
Question 7.2.
Given a positive real number Does there exist a positive real number and a positive integer such that for any -dimensional principally polarized abelian variety we have
2. Note that, except for Proposition 6.3, whose proof uses Nadel’s vanishing theorem, the results presented in this article hold over an algebraically closed field of any characteristic. We may then ask the following:
Question 7.3.
Let be a polarized abelian variety over a field of positive characteristic. Does the inequality still hold?
On the other hand, in [Mustata-Schwede] and in [Murayama], there are introduced the Frobenius-Seshadri constants of a line bundle at a point of a smooth variety over an algebraically closed field of positive characteristic and it is shown that the following inequalities hold:
where is the usual Seshadri constant at . This, together with our Theorem 1.1 3), suggests to study the following:
Question 7.4.
Let be an ample line bundle on an abelian variety defined over an algebraically closed field of positive characteristic. Compare the number and the -jets separation threshold where is the class of
This work is part of the author’s Phd thesis. The author would like to thank his advisor, Giuseppe Pareschi, for his invaluable guidance and help to improve both the exposition and the mathematical content of this article. The author also thanks to the anonymous referees for their comments, which were very useful to improve the presentation of this article and to clarify some proofs. This work was supported by the MIUR Excellence Department Project MatMod@TOV awarded to the Department of Mathematics of the University of Rome Tor Vergata.